# Magnetic disk winds in protoplanetary disks:

### v2

@@ 3.3 Magnetic disk winds and observational constraints
The second panel of Fig. 6 compares the co-evolution of the gas and dust radii. These radii obtained from ProDiMo continuum emissions are same as those presented in Fig. 4. The observations suggest that a PPD has its gas radius larger than the dust radius by up to a factor of 4, as indicated by the diagonal dashed lines in the figure. Similarly, the disk radii inferred from simulations consistently show $R_{\rm gas}>R_{\rm dust}$. As elaborated in Section 3.2, a smaller dust radius can be attributed to dust growth and drift and it is also related to a drop in the dust opacity coefficient. The models with high $\alpha_{\rm MRI}$, GV-2 and WI-2 show evolution towards top right, as a consequence of gravitoviscous spreading. The low $\alpha$ models with winds, WI-3 and WI-4, deviate significantly from the observed trends. As discussed in the context of Fig. 4, the overpowering magnitude of the winds in this case results in small disks, and increased dust drift results in an increasingly compact disk in Band 6 continuum. We note that the only difference between models WI-2, WI-3 and WI-4 is viscous $\alpha$. Thus, when the disk winds are involved, the exact magnitude of turbulence plays a significant role in overall disk evolution. The adjusted wind model WI-3a is much more consistent with the observational limits, as compared to WI-3. However, it still leaves some room for improvement, as the ratio of gas to dust radius is somewhat larger than $R_{\rm g}=4R_{\rm d}$ line. We note that the aim of model WI-3a is not to fit this wind model perfectly to the observations. Rather we adjust this model to show trends of varying the parameters associated with disk winds ($C_{\beta}$ and $C_{H}$) and demonstrate that a disk with reasonable characteristics can be obtained at low alpha without fine tuning.
@@
In Table 4, we list some of the global disk quantities that characterize the disks at 1 Myr. The numbers reinforce several conclusions made earlier and here we discuss some of the key points in terms of the mass budget of the PPDs. Model GV-4 produces a massive disk capable of forming envelopes of hundreds of gas giants, while WI-4 and WI-3 do not have enough mass to form a single gas giant. Most of the dust mass is in the grown dust, which is expected from dust dynamics occurring in a PPD. The low viscosity increases the fragmentation barrier for dust grains and the disk winds increase the inward dust drift, resulting in much more grown dust mass as compared to the small dust in these models. The global bulk dust to gas ratio ($\zeta_{\rm d2g}=M_{\rm disk,dust}/M_{\rm disk,gas}$) is approximately between $0.1-5\%$ and in most cases, the dust is depleted below the canonical ISM value of $1\%$. The “efficiency” of star formation can be inferred from the stellar mass, as all models start from the same mass reservoir of 0.83 $M_{\odot}$. With gravitoviscous evolution, the lower $\alpha$ models produce a star with lower mass, as the diminished angular momentum transport maintains a relatively massive disk. With the inclusion of winds, however, we see that the efficiency of star formation is almost independent of disk $\alpha$ and varies with the strength of the winds. The total mass lost to the winds is inversely proportional to viscous $\alpha$, as these disks produce stronger winds via its dependence on the gas surface density. However, a larger amount of net mass lost to the winds does not necessarily mean a less massive disk. E.g., in the case of WI-3a, the wind stress is attenuated by increasing $C_{\beta}$, while the wind mass loss rate is increased by increasing $C_{H}$. Comparing against WI-3, the disk mass is larger, while the mass lost to the winds comes at the cost of the final mass of the star. Finally, the $t_{\rm disk}=M_{\rm gas}$/$\dot{M}_{\star}$, where $\dot{M}_{\star}$ is stellar accretion rate, is calculated at 1Myr. This rough estimate of the disk lifetime implies that the disks with magnetic winds evolve faster with a noticeably shorter lifespan.
Table 4: Comparison of global disk properties at 1 Myr
@@ 3.4 Theoretical expectations from disk winds
In a recent review on protoplanetary disk demographics, Manara et al. (2023b) illustrate some potential population level differences that can distinguish disks that evolve via predominantly viscous processes versus those driven primarily by disk winds. This difference arises because the action of magnetic winds tends to shrink the disk size as opposed to its viscous spreading, and it is best distinguished via the divergence of the evolutionary tracks in the mass-radius plane. Fig. 7 depicts such a plot for the simulated models. Here the solid lines depict the theoretical centrifugal radius plotted against the disk mass, the latter corresponding to the gas mass enclosed within the centrifugal disk. $R_{\rm cf}$ is smoothed in time with a moving boxcar average with a box size of 0.05 Myr, hence these quantities correspond to the first rows of Figs. 4 and 5. The dashed lines show the disk sizes in 12CO, obtained from synthetic emissions and plotted against the synthetic disk mass at that time. The latter is calculated as 100 times the dust mass derived from the continuum emissions, since many studies still use this approximation. The diamonds mark endpoints of the simulations at 1 Myr. The scatter plot corresponds to the observed disks in Lupus cluster as seen in the ALMA survey data, where the disk masses are obtained from the Band 6 dust content (Ansdell et al., 2016; Miotello et al., 2023). The region on the extreme right represents the early and embedded GI-prone state when the disk is massive, while the trajectory will eventually move left towards lower mass and eventual dispersal, presumably via photoevaporation. The larger observed radius as compared to $R_{\rm cf}$ is seen in almost all models, which may be an artifact of a 90% flux threshold or additionally, of the criterion of surface density threshold in calculating $R_{\rm cf}$. On the other hand, the observed disk masses are systematically smaller, which are a result of significant dust depletion as well as growth during the disk evolution.
@@
Figure 7: Evolution of the disk models in the mass-radius plane, showing how the wind models may diverge from the gravitoviscous models after the disk achieves its maximum mass. The solid lines depict theoretical disk size, $R_{\rm cf}$, versus gas mass within the centrifugal disk. The dashed lines show $R_{\rm gas,^{12}CO}$ obtained from synthetic emissions versus disk mass, the latter calculated as 100 times the dust mass obtained from the synthetic continuum emission. The diamonds mark the end points at 1 Myr and the blue scatter plot shows observed values from an ALMA survey of the Lupus cluster (Ansdell et al., 2016; Miotello et al., 2023).
@@
The gravitoviscous models without disk winds spread monotonically, and end up with either a large disk radius (GV-2) or a large disk mass (GV-4). The low viscosity GV-4 model is able to hide most of its gas mass, however, such massive disks which are a significant fraction of the stellar mass typically exhibit GI spirals. Thus, GV models may be congruent with the observations of the largest or the most massive of the Class II disks. When both gravitoviscous and wind-driven accretion are present, the disk spreading is reduced and the radius evolution is essentially a result of the tug-of-war between these two processes. Hence, with the inclusion of disk winds, PPDs tend to evolve towards the diagonal in Fig. 7, so that they end up smaller and less massive. The disk wind models with low alpha, WI-3 and WI-4, produce compact disks that lie outside of observational expectations, although since $R_{\rm gas,^{12}CO}>R_{\rm cf}$, model WI-3 moves towards the observed values. Model WI-2 with a larger value of $\alpha$ as well as WI-3a are more or less congruent with the observed ranges. $R_{\rm cf}$ in WI-2 contracts near the end of its evolution, implying the size does not necessarily increase monotonically when disk winds are considered, although the synthetic disk size in $R_{\rm gas,^{12}CO}$ increases with time. These results are qualitatively consistent with the expectations from previous models of disk winds, and imply that magnetic disk winds may be necessary to explain observed distributions (e.g., Bai et al., 2016; Tabone et al., 2022). Fig. 7 demonstrates that in reality, both the measurements of disk sizes and masses contain several assumptions and significant uncertainties to draw decisive conclusions about the disk driving mechanism.
@@
At this point, we discuss the dimensionless quantity of magnetic lever arm, $\lambda$, which is typically assumed as a parameter for wind-driven accretion models that are analogous to simple viscous $\alpha$-disks (Kimmig et al., 2020; Alessi & Pudritz, 2022; Tabone et al., 2022). This parameter characterizes the removal of angular momentum and it is essentially the ratio of specific angular momentum evacuated by the wind to that in the underlying Keplerian region that launches the wind. In the classical picture of magnetocentrifugal winds, the magnetic field lines rotate rigidly inside of the Alfvén surface, so that the gas leaving as a wind increases its specific angular momentum (Blandford & Payne, 1982b). If a field line anchored at a radial distance $r_{0}$ crosses the Alfv́en surface at $r_{A}$, the magnetic lever arm can be written as
@@
In the simulations, the local values of $\lambda$ fluctuate severely. However, a global $\lambda$ can be calculated, which is averaged over the entire centrifugal disk at a given time. Fig. 8 shows the evolution of this inferred $\lambda$, which is smoothed with moving boxcar average with a box length of 0.05 Myr. For all models, $\lambda$ decreases initially, especially during the Class I phase, while its value remains remarkably constant at later times. During Class 0 and I stages, the disk is massive and also relatively compact, which results in large average gas surface densities. This implies greater wind torques at early times and a correspondingly large values of $\lambda$. For GV-2 and GV-4, where disk winds are not included self-consistently, $\lambda$ is relatively large at later times as well. This may be a result of a larger average $\Sigma_{\rm g}$, as disk winds are not included self-consistently in the GV models. For the disk wind models, the lever arms converge to $\lambda\approx 1-2$, which is in excellent agreement with both expectations from MHD simulations and observational constraints. The global values of $\lambda$ are relatively steady during Class II stage in all cases. This suggests that simpler disk wind models that start with an MMSN disk and adopt a constant lever arm may also be suitable for long-term evolution.
@@ 4 Discussions and conclusions
Magnetic disk winds play a crucial role in the evolution of protoplanetary disks, as suggested by both observational evidence and theoretical insights over the last two decades. In this study, we introduce a simple, semi-analytical formulation of magnetic disk winds in the thin-disk limit. The results of local shearing box simulations can be stitched together to obtain analytical radial dependence of the wind properties throughout the disk (Bai, 2013). These fitting formulae assume a power-law dependence of wind mass loss rate and surface stress as a function of local disk properties. We improve upon these results and take into account several corrections, e.g., due to changing stellar mass and FUV luminosity. We incorporate this model of magnetic disk winds self-consistently in FEOSAD code, which simulates the formation of protoplanetary disks and their evolution over long term.
@@
The wind mass loss rate and torque result in a sink term in mass continuity and momentum equations, respectively, while acting over the extent of the centrifugal disk. The corrections to other relevant equations such as energy conservation and those for small dust are also considered. The two free parameters in the disk wind model are $C_{\beta}$, which arises from the ideal MHD approximation used for the disk evolution, and $C_{\rm H}$, which originates in the relatively unconstrained mass loss rate as predicted by the shearing box simulations. The simulation results depend sensitively on the factor $C_{\beta}$, which scales the disk plasma $\beta_{0}$ in our ideal MHD models. However, we note that this parameter is not arbitrary, as it is chosen so that the quantity $\beta_{0}\times C_{\beta}$ over majority of the disk is of the order of $10^{4}$. This expected value is consistent with $\beta_{0}$ in protoplanetary disks from both observations and simulations, although these estimates vary by orders of magnitude. If our model is correct, only a narrow range of $C_{\beta}$ may be suitable for disk evolution, implying that there could be an underlying mechanism – perhaps some self-regulation process during magnetic transport – that causes different protoplanetary disks to have similar magnetic $\beta_{0}$ values. Moreover, $C_{\beta}$ may not be constant in space and time, and this uncertainty makes it challenging to compare models against observations. For the model to work without $C_{\beta}$, solving the non-ideal MHD equations is necessary. An alternative approach could be capping $B_{z}$ to a reasonable upper limit corresponding to observational constraints, e.g, from paleomagnetism of meteorites in the early solar system (Fu et al., 2014; Borlina et al., 2021). This approximation handles the two limiting cases of no and strong diffusivity correctly, as in the limit of strong magnetic diffusion $B_{z}$ will be approximately constant. Despite the aforementioned drawbacks, our model gives us a reasonable insight into the long-term evolution of wind-driven disks.
@@
Previously, FEOSAD considered gravitoviscous evolution, wherein a PPD forms from the collapsing parent cloud and subsequently evolves via combined action of turbulent viscosity and gravitational torques. In this framework, we now conduct numerical experiments, wherein the effects from disk winds are also included. In our limited parameter space study, three values of viscous $\alpha$ were considered ($0.01$, $10^{-3}$ and $10^{-4}$). Synthetic observations were obtained by post-processing the simulation results with radiation thermo-chemical code ProDiMo. For comparison with observations, we aggregated available ALMA survey data for nearby star forming regions. Here we validated our magnetic disk wind model with respect to observational constraints, while the disk properties and characteristics of the magnetic winds will be investigated in detail in a subsequent paper. Without discussing the individual models in detail, our salient findings can be summarized as follows.
@@
    PPDs evolving under gravitoviscous torques only (without disk winds), typically result in Class II disks that are too large in mass, and often in size, in comparison with observational expectations.
@@
    With the inclusion of magnetic disk winds, the resulting disks are smaller and less massive. This is expected, as advective evolution caused by vertical angular momentum transport from disk winds tends to suppress the outwards gravitoviscous spread and promote inward mass transport.
@@
    Some of the models with winds produce Class II disks that are unreasonably small, in terms of dust disk size as well as the true disk mass. Obtaining a reasonable fit to the observations is possible with a factor of few adjustment in the two disk wind model parameters $C_{\beta}$ and $C_{\rm H}$.
@@
    The synthetic observations of global disk properties - gas radius in 12CO, dust radius in Band 6 continuum, dust content as calculated by Band 6 flux as well as 13CO and C18O line fluxes - move in the direction of better agreement with the ALMA survey data, when disk winds are included.
@@
    Similar to previous gravitoviscous simulations, models including winds also show formation of a dead zone, pressure maxima, and dusty rings with enhanced dust-to-gas ratio in the inner disk region. These can assist rapid planetesimal formation, and point towards planet formation starting early in the Class I stage.
@@
    The divergence of evolutionary tracks with respect to disk winds may be seen on the disk mass vs radius plane, where careful choice of disk wind parameters leads to models that are more consistent with the observed ranges.
@@
    The ratio of synthetic gas to dust radii stays within the observed ranges of 1-4 for the gravitoviscous models. At low viscous $\alpha$, the models which include disk winds may produce anomalously small dust disks, due to the loss of small dust and efficient migration of grown dust. By changing model parameters, the magnetic winds can be attenuated so that the resulting disk fits the observational constraints.
@@
    The synthetic spectral indices for all models are broadly consistent with the survey data throughout their evolution. As a disk evolves, an increasing $\alpha_{\rm mm}$ indicates grain growth and decrease in Band 6 flux indicates decrease in the dust content.
@@
    In agreement with both simulations and observations, the inferred effective magnetic lever arm for an entire disk is between $\approx 1-3$ for the models that include action of disk winds, and it stays remarkably steady during the Class II stage.
In conclusion, our magnetic disk wind model is able to reproduce the observational properties, while simultaneously satisfying theoretical expectations over the long term ($\sim 1$ Myr), wherein the disk forms and evolves through the Class 0, I, and II stages. Even now, most models of planet formation start with a fully formed MMSN protoplanetary disk and consider only its viscous evolution. Recent observations of PPDs put severe constraints on planet formation, as Class II disks do not contain enough dust mass to support the observed population of exoplanets (Greaves & Rice, 2010; Tychoniec et al., 2020). This suggests that planetesimals and planetary embryos start forming earlier in the embedded stages, when the gravitational torques are significant, while disk winds control the fundamental disk properties such as its size and the mass budget, affecting their subsequent evolution. Our model marks a step forward in the direction of representing a more complete picture of such disks, wherein the concurrent effects from viscous, gravitational, and magnetic disk wind processes guide the disk evolution and ultimately decide the fate of the planetary system that emerges.

### v3

@@ 3.3 Magnetic disk winds and observational constraints
The second panel of Fig. 6 compares the co-evolution of the gas and dust radii. These radii obtained from ProDiMo continuum emissions are same as those presented in Fig. 4. The observations suggest that a PPD has its gas radius larger than the dust radius by up to a factor of 4, as indicated by the diagonal dashed lines in the figure. Similarly, the disk radii inferred from simulations consistently show $R_{\rm gas}>R_{\rm dust}$. As elaborated in Sect. 3.2, a smaller dust radius can be attributed to dust growth and drift and it is also related to a drop in the dust opacity coefficient. The models with high $\alpha_{\rm MRI}$, GV-2 and WI-2 show evolution toward top right, as a consequence of gravitoviscous spreading. The low $\alpha$ models with winds, WI-3 and WI-4, deviate significantly from the observed trends. As discussed in the context of Fig. 4, the overpowering magnitude of the winds in this case results in small disks, and increased dust drift results in an increasingly compact disk in Band 6 continuum. We note that the only difference between models WI-2, WI-3 and WI-4 is viscous $\alpha$. Thus, when the disk winds are involved, the exact magnitude of turbulence plays a significant role in overall disk evolution. The adjusted wind model WI-3a is much more consistent with the observational limits, as compared to WI-3. However, it still leaves some room for improvement, as the ratio of gas to dust radius is somewhat larger than $R_{\rm g}=4R_{\rm d}$ line. We note that the aim of model WI-3a is not to fit this wind model perfectly to the observations. Rather we adjust this model to show trends of varying the parameters associated with disk winds ($C_{\beta}$ and $C_{H}$) and demonstrate that a disk with reasonable characteristics can be obtained at low alpha without fine tuning.
@@
In Table 4, we list some of the global disk quantities that characterize the disks at one Myr. The numbers reinforce several conclusions made earlier and here we discuss some of the key points in terms of the mass budget of the PPDs. Model GV-4 produces a massive disk capable of forming envelopes of hundreds of gas giants, while WI-4 and WI-3 do not have enough mass to form a single gas giant. Most of the dust mass is in the grown dust, which is expected from dust dynamics occurring in a PPD. The low viscosity increases the fragmentation barrier for dust grains and the disk winds increase the inward dust drift, resulting in much more grown dust mass as compared to the small dust in these models. The global bulk dust to gas ratio ($\zeta_{\rm d2g}=M_{\rm disk,dust}/M_{\rm disk,gas}$) is approximately between $0.1-5\%$ and in most cases, the dust is depleted below the canonical ISM value of $1\%$. The “efficiency” of star formation can be inferred from the stellar mass, as all models start from the same mass reservoir of 0.83 $M_{\odot}$. With gravitoviscous evolution, the lower $\alpha$ models produce a star with lower mass, as the diminished angular momentum transport maintains a relatively massive disk. With the inclusion of winds, however, we see that the efficiency of star formation is almost independent of disk $\alpha$ and varies with the strength of the winds. The total mass lost to the winds is inversely proportional to viscous $\alpha$, as these disks produce stronger winds via its dependence on the gas surface density. However, a larger amount of net mass lost to the winds does not necessarily mean a less massive disk. For example, in the case of WI-3a, the wind stress is attenuated by increasing $C_{\beta}$, while the wind mass loss rate is increased by increasing $C_{H}$. Comparing against WI-3, the disk mass is larger, while the mass lost to the winds comes at the cost of the final mass of the star. Finally, the $t_{\rm disk}=M_{\rm gas}$/$\dot{M}_{\star}$, where $\dot{M}_{\star}$ is stellar accretion rate, is calculated at one Myr. This rough estimate of the disk lifetime implies that the disks with magnetic winds evolve faster with a noticeably shorter lifespan.
Table 4: Comparison of global disk properties at one Myr.
@@ 3.4 Theoretical expectations from disk winds
In a recent review on protoplanetary disk demographics, Manara et al. (2023b) illustrate some potential population level differences that can distinguish disks that evolve via predominantly viscous processes versus those driven primarily by disk winds. This difference arises because the action of magnetic winds tends to shrink the disk size as opposed to its viscous spreading, and it is best distinguished via the divergence of the evolutionary tracks in the mass-radius plane. Figure 7 depicts such a plot for the simulated models. Here the solid lines depict the theoretical centrifugal radius plotted against the disk mass, the latter corresponding to the gas mass enclosed within the centrifugal disk. $R_{\rm cf}$ is smoothed in time with a moving boxcar average with a box size of 0.05 Myr, hence these quantities correspond to the first rows of Figs. 4 and 5. The dashed lines show the disk sizes in 12CO, obtained from synthetic emissions and plotted against the synthetic disk mass at that time. The latter is calculated as 100 times the dust mass derived from the continuum emissions, since many studies still use this approximation. The diamonds mark endpoints of the simulations at one Myr. The scatter plot corresponds to the observed disks in Lupus cluster as seen in the ALMA survey data, where the disk masses are obtained from the Band 6 dust content (Ansdell et al., 2016; Miotello et al., 2023). The region on the extreme right represents the early and embedded GI-prone state when the disk is massive, while the trajectory will eventually move left toward lower mass and eventual dispersal, presumably via photoevaporation. The larger observed radius as compared to $R_{\rm cf}$ is seen in almost all models, which may be an artifact of a 90% flux threshold or additionally, of the criterion of surface density threshold in calculating $R_{\rm cf}$. On the other hand, the observed disk masses are systematically smaller, which are a result of significant dust depletion as well as growth during the disk evolution.
@@
Figure 7: Evolution of the disk models in the mass-radius plane showing how the wind models may diverge from the gravitoviscous models after the disk achieves its maximum mass. The solid lines depict theoretical disk size, $R_{\rm cf}$, versus gas mass within the centrifugal disk. The dashed lines show $R_{\rm gas,^{12}CO}$ obtained from synthetic emissions versus disk mass, the latter calculated as 100 times the dust mass obtained from the synthetic continuum emission. The diamonds mark the end points at one Myr and the blue scatter plot shows observed values from an ALMA survey of the Lupus cluster (Ansdell et al., 2016; Miotello et al., 2023).
@@
The gravitoviscous models without disk winds spread monotonically, and end up with either a large disk radius (GV-2) or a large disk mass (GV-4). The low viscosity GV-4 model is able to hide most of its gas mass; however, such massive disks, which are a significant fraction of the stellar mass, typically exhibit GI spirals. Thus, GV models may be congruent with the observations of the largest or the most massive of the Class II disks. When both gravitoviscous and wind-driven accretion are present, the disk spreading is reduced and the radius evolution is essentially a result of the tug-of-war between these two processes. Hence, with the inclusion of disk winds, PPDs tend to evolve toward the diagonal in Fig. 7 so that they end up smaller and less massive. The disk wind models with low alpha, WI-3 and WI-4, produce compact disks that lie outside of observational expectations, although since $R_{\rm gas,^{12}CO}>R_{\rm cf}$, model WI-3 moves toward the observed values. Model WI-2 with a larger value of $\alpha$ as well as WI-3a are more or less congruent with the observed ranges. $R_{\rm cf}$ in WI-2 contracts near the end of its evolution, implying the size does not necessarily increase monotonically when disk winds are considered, although the synthetic disk size in $R_{\rm gas,^{12}CO}$ increases with time. These results are qualitatively consistent with the expectations from previous models of disk winds, and imply that magnetic disk winds may be necessary to explain observed distributions (e.g., Bai et al., 2016; Tabone et al., 2022). Figure 7 demonstrates that in reality, both the measurements of disk sizes and masses contain several assumptions and significant uncertainties to draw decisive conclusions about the disk driving mechanism.
@@
At this point, we discuss the dimensionless quantity of magnetic lever arm, $\lambda$, which is typically assumed as a parameter for wind-driven accretion models that are analogous to simple viscous $\alpha$-disks (Kimmig et al., 2020; Alessi & Pudritz, 2022; Tabone et al., 2022). This parameter characterizes the removal of angular momentum and it is essentially the ratio of specific angular momentum evacuated by the wind to that in the underlying Keplerian region that launches the wind. In the classical picture of magnetocentrifugal winds, the magnetic field lines rotate rigidly inside of the Alfvén surface so that the gas leaving as a wind increases its specific angular momentum (Blandford & Payne, 1982b). If a field line anchored at a radial distance $r_{0}$ crosses the Alfv́en surface at $r_{A}$, the magnetic lever arm can be written as
@@
In the simulations, the local values of $\lambda$ fluctuate severely. However, a global $\lambda$ can be calculated, which is averaged over the entire centrifugal disk at a given time. Figure 8 shows the evolution of this inferred $\lambda$, which is smoothed with moving boxcar average with a box length of 0.05 Myr. For all models, $\lambda$ decreases initially, especially during the Class I phase, while its value remains remarkably constant at later times. During Class 0 and I stages, the disk is massive and also relatively compact, which results in large average gas surface densities. This implies greater wind torques at early times and a correspondingly large values of $\lambda$. For GV-2 and GV-4, where disk winds are not included self-consistently, $\lambda$ is relatively large at later times as well. This may be a result of a larger average $\Sigma_{\rm g}$, as disk winds are not included self-consistently in the GV models. For the disk wind models, the lever arms converge to $\lambda\approx 1-2$, which is in excellent agreement with both expectations from MHD simulations and observational constraints. The global values of $\lambda$ are relatively steady during Class II stage in all cases. This suggests that simpler disk wind models that start with an MMSN disk and adopt a constant lever arm may also be suitable for long-term evolution.
@@ 4 Discussions and conclusions
Magnetic disk winds play a crucial role in the evolution of protoplanetary disks, as suggested by both observational evidence and theoretical insights over the past two decades. In this study, we have introduced a simple semi-analytical formulation of magnetic disk winds in the thin-disk limit. The disk wind model is based on the insights gained from local shearing box simulations. The results of such local simulations can be stitched together to obtain analytical radial dependence of the wind properties throughout the disk (Bai, 2013). These fitting formulae assume a power-law dependence of wind mass loss rate and surface stress as a function of local disk properties. We improved upon these results and took into account several corrections, for example, due to changing stellar mass and FUV luminosity. We incorporated this model of magnetic disk winds self-consistently in the FEOSAD code, which simulates the formation of protoplanetary disks and their evolution over the long term.
@@
The wind mass loss rate and torque result in a sink term in mass continuity and momentum equations, respectively, while acting over the extent of the centrifugal disk. The corrections to other relevant equations, such as energy conservation and those for small dust, were also considered. The two free parameters in the disk wind model are $C_{\beta}$, which arises from the ideal MHD approximation used for the disk evolution, and $C_{\rm H}$, which originates in the relatively unconstrained mass loss rate as predicted by the shearing box simulations. The simulation results depend sensitively on the factor $C_{\beta}$, which scales the disk plasma $\beta_{0}$ in our ideal MHD models. However, we note that this parameter is not arbitrary, as it is chosen so that the quantity $\beta_{0}\times C_{\beta}$ over majority of the disk is of the order of $10^{4}$. This expected value is consistent with $\beta_{0}$ in protoplanetary disks from both observations and simulations, although these estimates vary by orders of magnitude. If our model is correct, only a narrow range of $C_{\beta}$ may be suitable for disk evolution, implying that there could be an underlying mechanism – perhaps some self-regulation process during magnetic transport – that causes different protoplanetary disks to have similar magnetic $\beta_{0}$ values. Moreover, $C_{\beta}$ may not be constant in space and time, and this uncertainty makes it challenging to compare models against observations. For the model to work without $C_{\beta}$, solving the non-ideal MHD equations is necessary. An alternative approach could be capping $B_{z}$ to a reasonable upper limit corresponding to observational constraints, for example, from paleomagnetism of meteorites in the early Solar System (Fu et al., 2014; Borlina et al., 2021). This approximation handles the two limiting cases of no and strong diffusivity correctly, as $B_{z}$ will be approximately constant in the limit of strong magnetic diffusion. Despite the aforementioned drawbacks, our model gives us a reasonable insight into the long-term evolution of wind-driven disks.
@@
Previously, FEOSAD considered gravitoviscous evolution, wherein a PPD forms from the collapsing parent cloud and subsequently evolves via combined action of turbulent viscosity and gravitational torques. In this framework, we conducted numerical experiments that incorporated the effects of disk winds. In our limited parameter space study, three values of viscous $\alpha$ were considered ($0.01$, $10^{-3}$, and $10^{-4}$). Synthetic observations were obtained by post-processing the simulation results with the radiation thermo-chemical code ProDiMo. For comparison with observations, we aggregated available ALMA survey data for nearby star forming regions. We validated our magnetic disk wind model with respect to observational constraints, while the disk properties and characteristics of the magnetic winds will be investigated in detail in a subsequent paper. Without discussing the individual models in detail, our salient findings can be summarized as follows:
@@
    Protoplanetary disks evolving under gravitoviscous torques only (without disk winds) typically result in Class II disks that are too large in mass, and often in size, in comparison with observational expectations.
@@
    With the inclusion of magnetic disk winds, the resulting disks are smaller and less massive. This is expected, as advective evolution caused by vertical angular momentum transport from disk winds tends to suppress the outward gravitoviscous spread and promote inward mass transport.
@@
    Some of the models with winds produce Class II disks that are unreasonably small in terms of dust disk size as well as the true disk mass. Obtaining a reasonable fit to the observations is possible with a few adjustments in the two disk wind model parameters $C_{\beta}$ and $C_{\rm H}$.
@@
    The synthetic observations of global disk properties - gas radius in 12CO, dust radius in Band 6 continuum, dust content as calculated by Band 6 flux, and 13CO and C18O line fluxes - move in the direction of better agreement with the ALMA survey data when disk winds are included.
@@
    Similar to previous gravitoviscous simulations, models including winds also show the formation of a dead zone, pressure maxima, and dusty rings with an enhanced dust-to-gas ratio in the inner disk region. These can assist rapid planetesimal formation and point toward planet formation starting early in the Class I stage.
@@
    The divergence of evolutionary tracks with respect to disk winds may be seen on the disk mass versus radius plane, where careful choice of disk wind parameters leads to models that are more consistent with the observed ranges.
@@
    The ratio of synthetic gas to dust radii stays within the observed ranges of one to four for the gravitoviscous models. At a low viscous $\alpha$, the models that include disk winds may produce anomalously small dust disks due to the loss of small dust and efficient migration of grown dust. By changing model parameters, the magnetic winds can be attenuated so that the resulting disk fits the observational constraints.
@@
    The synthetic spectral indices for all models are broadly consistent with the survey data throughout their evolution. As a disk evolves, an increasing $\alpha_{\rm mm}$ indicates grain growth, and a decrease in Band 6 flux indicates a decrease in the dust content.
@@
    In agreement with simulations and observations, the inferred effective magnetic lever arm for an entire disk is between approximately one and three for the models that include the action of disk winds, and it stays remarkably steady during the Class II stage.
In conclusion, our magnetic disk wind model is able to reproduce the observational properties while simultaneously satisfying theoretical expectations over the long term (over one million years), wherein the disk forms and evolves through the Class 0, I, and II stages. Even now, most models of planet formation start with a fully formed MMSN protoplanetary disk and consider only its viscous evolution. Recent observations of PPDs put severe constraints on planet formation, as Class II disks do not contain enough dust mass to support the observed population of exoplanets (Greaves & Rice, 2010; Tychoniec et al., 2020). This suggests that planetesimals and planetary embryos start forming earlier in the embedded stages, when the gravitational torques are significant, while disk winds control the fundamental disk properties, such as its size and the mass budget, affecting their subsequent evolution. Our model marks a step forward in the direction of representing a more complete picture of such disks, wherein the concurrent effects from viscous, gravitational, and magnetic disk wind processes guide the disk evolution and ultimately decide the fate of the planetary system that emerges.